# Neighbour-Driven Gaussian Process Variational Autoencoders for Scalable Structured Latent Modelling

### v2

@@ D.2.2 Markovian GPVAEs
[…]
@@
[…]
@@
[…]
@@ D.3 SGPBAE
is a neural network $p(\mathbf{Y}\mid\mathbf{Z},\boldsymbol{\theta})$ with weights/biases $\boldsymbol{\theta}$ regarded as random variables  [^4] To remain notation consistency. The paper uses $\boldsymbol{\varphi}$ to represent the decoder’s parameters and $\boldsymbol{\theta}$ for kernel parameters instead.. Therefore, defining $\boldsymbol{\Psi}=\{\psi,\mathbf{S},\mathbf{U},\boldsymbol{\theta}\}$, the potential energy function is given by
@@
$\begin{cases}\Delta\boldsymbol{\Theta}&=\eta\mathbf{M}^{-1}\mathbf{r},\\ \Delta\mathbf{r}&=-\eta\nabla\tilde{U}(\boldsymbol{\Theta})-\eta\mathbf{C}\mathbf{M}^{-1}\mathbf{r}+\mathcal{N}(\mathbf{0},2\eta(\mathbf{C}-\tilde{\mathbf{B}})),\end{cases}$
@@
[…]
Prediction
They collect $K$ samples $\{\mathbf{\Psi}_{k}\}_{k=1}^{K}=\{\psi_{k},\mathbf{S}_{k},\mathbf{U}_{k},\mathbf{\theta}_{k}\}_{k=1}^{K}$ from the MCMC sampling. For each sample $\mathbf{\Psi}_{k}$, the prediction for a newly introduced location $\mathbf{x}_{*}$ is computed as follows:
  • $q(\mathbf{z}_{*}^{k}\mid\mathbf{u}_{k},\mathbf{S}_{k},\psi_{k})=\mathcal{N}\left(\mathbf{z}_{*}^{k}\mid\mathbf{K}_{\mathbf{x}_{*}\mathbf{S}_{k}\mid\psi_{k}}\mathbf{K}_{\mathbf{S}_{k}\mathbf{S}_{k}\mid\psi_{k}}^{-1}\mathbf{u}_{k},\operatorname{diag}\left[\mathbf{K}_{\mathbf{x}_{*}\mathbf{x}_{*}\mid\psi_{k}}-\mathbf{K}_{\mathbf{x}_{*}\mathbf{S}\mid\psi_{k}}\mathbf{K}_{\mathbf{S}\mathbf{S}\mid\psi_{k}}^{-1}\mathbf{K}_{\mathbf{S}\mathbf{x}_{*}}\right]+\sigma^{2}_{\psi_{k}}\right)$.
  • Sample $\mathbf{z}_{*}^{k}$ from $q(\mathbf{z}_{*}^{k})$ and generate $\mathbf{y}_{*}\sim p_{\mathbf{\theta}_{k}}(\mathbf{y}_{*}\mid\mathbf{z}_{*}^{k})$.

### v3

@@ D.2.2 Markovian GPVAEs
Computing the state-space posterior distribution at a newly introduced time point $t_{*}$ relies on the smoothed distributions obtained at the adjacent time points. More concretely, let $t_{-}$ and $t_{+}$ denote the immediate predecessor and successor time points, respectively, and the predictive $q(\mathbf{s}_{t_{*}})$ can be obtained as $q(\mathbf{s}_{t_{*}})=\int p(\mathbf{s}_{t_{*}}\mid\mathbf{s}_{t_{-}},\mathbf{s}_{t_{+}})q(\mathbf{s}_{t_{-}},\mathbf{s}_{t_{+}})d\mathbf{s}_{t_{-}}d\mathbf{s}_{t_{+}}$. More details can be found in Appendix A.1 of Adam et al. (2020). Using the relationship $\mathbf{z}_{t_{*}}=\mathbf{H}\mathbf{s}_{t_{*}}$, we can obtain the predictive posterior $q(\mathbf{z}_{t_{*}})$ and $q(\mathbf{y}_{*})$.
@@
$=\mathbf{Q}_{-*}\mathbf{A}_{*+}^{\top}\mathbf{Q}_{*+}^{-1}-\mathbf{Q}_{-*}\mathbf{A}_{*+}^{\top}\mathbf{Q}_{-+}^{-1}(\mathbf{Q}_{-+}-\mathbf{Q}_{*+})\mathbf{Q}_{*+}^{-1}\quad\text{(Woodbury identity)}$
@@
$q(\mathbf{s}_{t_{*}})=\int\mathcal{N}(\mathbf{s}_{t_{*}}\mid\mathbf{P}\mathbf{v},\mathbf{T})q(\mathbf{v})d\mathbf{v}=\mathcal{N}\left(\mathbf{s}_{t_{*}}\mid\mathbf{P}\begin{bmatrix}\mathbf{m}_{t_{-}}^{s}\\ \mathbf{m}_{t_{+}}^{s}\end{bmatrix},\mathbf{P}\mathbf{\Sigma_{-+}}\mathbf{P}^{\top}+\mathbf{T}\right).$
@@ D.3 SGPBAE
is a neural network $p(\mathbf{Y}\mid\mathbf{Z},\boldsymbol{\theta})$ with weights/biases $\boldsymbol{\theta}$ regarded as random variables  [^3] To remain notation consistency. The original paper uses $\boldsymbol{\varphi}$ to represent the decoder’s parameters and $\boldsymbol{\theta}$ for kernel parameters instead.. Therefore, defining $\boldsymbol{\Psi}=\{\psi,\mathbf{S},\mathbf{U},\boldsymbol{\theta}\}$, the potential energy function is given by
@@
[…]